a boat completes a trip from riverton to clear creek and back each weekday.the distance one way is 60 miles ,and the speed of the current in the river is 4 miles per hour .if the round trip takes 8 hours ,determine the speed of the boat in still water
Answer by lwsshak3(11628)

a boat completes a trip from riverton to clear creek and back each weekday.the distance one way is 60 miles ,and the speed of the current in the river is 4 miles per hour .if the round trip takes 8 hours ,determine the speed of the boat in still water
**
let x=speed of boat in still water
x+4=speed of boat with current
x-4=speed of boat against current
Travel time=distance/speed
..
60/(x+4)+60/(x-4)=8
LCD: (x+4)(x-4)
60(x-4)+60(x+4)=8(x+4)(x-4)=8(x^2-16)
60x-240+60x+240=8x^2-128
120x=8x^2-128
8x^2-120x-128=0
divide by 8
x^2-15x-16=0
(x+1)(x-16)=0
x=-1 (reject, x>0)
or
x=16
ans:
speed of boat in still water=16 mph